OpenStudy (anonymous):

Express sin 3x in terms of sin x. PLEASE HELP!

zepdrix (zepdrix):

lol dat user name. So it's \(\Large\rm \sin3x\), yes? Not to be confused with \(\Large\rm \sin^3x\)

OpenStudy (anonymous):

yes its sin 3x

OpenStudy (anonymous):

my username OP

zepdrix (zepdrix):

I've only played a few games with Riven, still trying to get the hang of her >.< I main Elise in Jungle.

zepdrix (zepdrix):

So I guess we need to make use of our Angle Identity Formula for Sine.

zepdrix (zepdrix):

Angle Addition I meant to say*

OpenStudy (anonymous):

lol honorary professor of mathematics plays league

zepdrix (zepdrix):

\[\Large\rm \sin(3x)=\sin(2x+x)\]

zepdrix (zepdrix):

\[\Large\rm \sin(a+b)=\sin a \cos b +\sin b \cos a\]

zepdrix (zepdrix):

Understand how we can use this formula to start breaking things down?

OpenStudy (anonymous):

yeah i know this i just didnt know i forgot to apply it

OpenStudy (anonymous):

so that would be sin2xcosx+sinxcos2x

zepdrix (zepdrix):

Ok great. First step is done.

zepdrix (zepdrix):

From there to deal with the 2x angles, we'll apply our `Double Angle` Identities.

zepdrix (zepdrix):

\[\Large\rm \sin(2x)=2\sin x \cos x,\qquad\qquad \cos(2x)=1-2\sin^2x\]

OpenStudy (anonymous):

so that would be \[2 \sin x \cos ^{2}x +\sin x \left( 1-2\sin ^{2} x\right)\]

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Mmm ok. So now we have a cos^2x to deal with. Remember your Pythagorean Identity for sine and cosine?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\Large\rm \sin^2x+\cos^2x=1\qquad\to\qquad \cos^2x=1-\sin^2x\]

OpenStudy (anonymous):

so you plug that in

zepdrix (zepdrix):

\[\Large\rm 2\sin x \left(1-\sin^2x\right) +\sin x \left( 1-2\sin ^{2} x\right)\]Yes. And it shouldn't be to bad after that. Just need to distribute some stuff... and combine like terms.

OpenStudy (anonymous):

so that gives you\[3\sin x -4\sin ^{3}x\]

zepdrix (zepdrix):

Yay good job!

OpenStudy (anonymous):

is that the final answer?

zepdrix (zepdrix):

Yes.

zepdrix (zepdrix):

It's now it terms of sin(x), only x as the angle.

OpenStudy (anonymous):

YES THANK YOU THANK YOU THANK YOU

zepdrix (zepdrix):

lol np :3